# And yet it moves: why irrotational waves induce a mean Lagrangian drift

### v2

@@ Abstract
The particle trajectories in irrotational, incompressible and inviscid deep-water surface gravity waves are open, leading to a net drift in the direction of wave propagation commonly referred to as the Stokes Drift, which is responsible for catalysing surface wave-induced mixing in the ocean and transporting marine debris. A balance between phase-averaged momentum density, kinetic energy density and vorticity for irrotational, monochromatic and periodic two-dimensional water waves is derived by working directly within the Lagrangian reference frame, which tracks particle trajectories as a function of their labels and time. This balance should be expected as all three of these quantities are conserved following particles in this system. Vorticity in particular is always conserved along particles in two-dimensional inviscid flow, and as such even in its absence it is the value of the vorticity which fundamentally sets the drift, which in the Lagrangian frame is identified as the phase-averaged momentum density of the system. A relationship between the drift and the geometric mean water level of particles is found at the surface and applications for potential new ways of inferring drift are discussed. Finally, an example of an initially quiescent fluid driven by a wavelike pressure disturbance is considered, showing how the net momentum and energy from the surface disturbance transfer to the wave field, recognizing the source of the mean Lagrangian drift as the net momentum required to generate an irrotational surface wave by any conservative force.
@@ 1 Introduction
Deep water surface gravity waves are ubiquitous in the global oceans, and affect the transport of heat, momentum and mass both along and across the air-sea interface (Melville, 1996; Deike, 2022). One crucial property of irrotational deep water waves is that the particle trajectories are not closed leading to a net drift in the direction of wave propagation commonly referred to as Stokes Drift (Stokes, 1847). Formally, the Stokes drift is defined as the difference between the mean Lagrangian and mean Eulerian currents,
@@
where $\bm{u}_{L}$ and $\bm{u}_{E}$ represent the Lagrangian and Eulerian currents respectively, and the overline indicates a time mean in each reference frame over a wave period. One often neglected issue with this definition is the validity of taking the difference between two quantities in different reference frames with different dependent variables and more importantly different definitions of averaging as the Lagrangian and Eulerian periods are not equal (Longuet-Higgins, 1986). To avoid this confusion, we will instead use the term ‘wave-induced mean Lagrangian drift’ to refer to the mean Lagrangian velocity of fluid particles over the Lagrangian wave period. In this paper we restrict ourselves to vanishing external Eulerian currents, i.e. the fluid would be quiescent without waves.
The wave-induced mean Lagrangian drift modulates upper ocean currents, affects the transport of buoyant pollutants, plankton and marine debris, and enhances vertical mixing via Langmuir circulation (Craik & Leibovich, 1976; Belcher, 2012; DiBenedetto et al., 2018). There is also evidence that this drift, or mean Lagrangian momentum density, can help with the interpretation of many central geometric, kinematic and dynamic properties of surface waves (Pizzo et al., 2023). Despite the elapse of over 175 years since its discovery, there is still confusion regarding the origins and interpretation of the wave-induced mean flow for irrotational surface gravity waves. Most derivations, including that of Stokes (1847), calculate the magnitude and direction of the drift from an asymptotic integration of the kinematic condition relating the Eulerian and Lagrangian velocities, which simply states that at a fixed point in time and space, the Eulerian and Lagrangian velocities are equal
@@
since at a fixed time a particle’s location is coincident with a fixed point in space. Thus the particle trajectories within a wave, which are fundamentally Lagrangian quantities, are derived from the Eulerian velocity fields. When done in this way, the mean Lagrangian drift appears to simply fall out of the math, and physical explanations for its existence tend to come post-factum. Why, fundamentally, should progressive irrotational surface waves induce a mean motion of water? What sets its magnitude and direction? Lastly, how is the mean Lagrangian drift related to other quantities such as vorticity and energy density? Such questions are the primary aim of this paper.
In section 2, we introduce the governing equations and relevant conditions for solving irrotational, incompressible and inviscid two-dimensional deep-water surface gravity waves in the Lagrangian reference frame, which tracks the trajectories of individual fluid parcels as a function of labelling coordinates. In this frame, the wave-induced mean Lagrangian drift is explicitly written as the average velocity of fluid parcels, and is thus identified as the mean Lagrangian momentum density of the system, physically motivated when one recalls momentum being equivalent to mass flux. In section 3, through an investigation of these equations, we show how the mean momentum density, or equivalently the drift, is related to the vorticity and energy density of irrotational water waves. Despite the flow being completely irrotational, it is precisely this strict condition that kinematically mandates a sheared mean Lagrangian drift, and thus we emphasize that even in irrotational flow, it is the vorticity that sets the magnitude and direction of the drift. We then dynamically connect the mean momentum and kinetic energy densities, showing that all monochromatic, irrotational and progressive waves with nonzero kinetic energy require a nonzero mean Lagrangian drift. Finally, we highlight a connection between the mean potential and mean kinetic energy densities through the mean pressure using a Bernoulli equation in the Lagrangian frame first outlined by Pizzo et al. (2023).
To further explore the dynamic relationship between these variables, in section 4 we consider the momentum and energy budgets within the physically motivated Lagrangian reference frame. To show how the wave-induced mean Lagrangian drift emerges from an initially quiescent flow, we consider a simple example where a still surface is resonantly driven by a wavelike pressure forcing. Through the integral momentum budgets, we find that all the momentum transferred to the waves from the surface forcing goes into the mean momentum, or mean Lagrangian drift. The same is shown to be true for the total energy.
@@ 2 The Lagrangian description of water waves
Lagrangian quantities track evolution following fixed fluid particles. Thus a complete two-dimensional Lagrangian description of a fluid requires calculating particle trajectories $\bm{x}(\alpha,\beta,t)$ as a function of particle labelling coordinates $(\alpha,\beta)$ and time. These Lagrangian labels are fixed following a particle (i.e. $\dot{\alpha}=\dot{\beta}=0$) where dots represent a total time derivative, so that the total time derivative becomes a simple partial derivative in the Lagrangian frame. The particle trajectories $\bm{x}(\alpha,\beta,t)$ represent a general time-dependent coordinate transformation between label space $(\alpha,\beta)$ and physical space $(x,y)$, with a corresponding Jacobian determinant
@@
where subscripts indicate partial derivatives. The Jacobian also allows us to easily change variables of differentiation. For example, the two dimensional incompressiblity condition in the Eulerian frame is denoted as
@@
which can be mapped to the Lagrangian frame as
$0=u_{x}+v_{y}=\frac{\partial(u,y)}{\partial(x,y)}+\frac{\partial(x,v)}{\partial(x,y)}=\frac{1}{\mathcal{J}}\bigg{(}\frac{\partial(u,y)}{\partial(\alpha,\beta)}+\frac{\partial(x,v)}{\partial(\alpha,\beta)}\bigg{)}=\\ \frac{1}{\mathcal{J}}\bigg{(}\frac{\partial(\dot{x},y)}{\partial(\alpha,\beta)}+\frac{\partial(x,\dot{y})}{\partial(\alpha,\beta)}\bigg{)}=\frac{1}{\mathcal{J}}\frac{\partial}{\partial t}\frac{\partial(x,y)}{\partial(\alpha,\beta)}=\frac{1}{\mathcal{J}}\frac{\partial\mathcal{J}}{\partial t}=0\,.$ (5)
Therefore incompressible flow requires that the Jacobian $\mathcal{J}$ be time independent and not equal zero anywhere in the flow (i.e. no sign changes) so that the mapping remains invertible. We could have determined this condition without calculation by remembering that the Jacobian determines how infinitessimal areas are mapped between label space and physical space. A small collection of particles ${\rm d}\alpha\,{\rm d}\beta$ must enclose the same physical area $\mathcal{J}^{-1}{\rm d}x\,{\rm d}y$ for all time or else the flow would be allowed to compress.
@@
where $p$ represents the pressure, and $g$ the acceleration due to gravity. Note that in the Lagrangian frame the nonlinearities arise in the pressure gradient terms and not in the inertia terms, in contrast to the Eulerian reference frame.
While incompressibility provides a constraint on our mapping between label space and physical space, there is still tremendous freedom in how we label our particles; this is known as the particle relabelling symmetry (Salmon, 1988), and represents a gauge freedom of fluid mechanics. Just as in electromagnetism, this gauge can be conveniently chosen to simply computations, but we leave it general for now.
Here, our physical system amounts to solving equations (6), (7) for variables $(x,y,p)$ as functions of $(\alpha,\beta,t)$, subject to the incompressibility condition (5) for a given labelling gauge choice. To close the system, we impose boundary conditions at the free surface and the bottom. As part of our labelling freedom, we label particles at the surface with $\beta=0$, which makes the evaluation of surface quantities straightforward. This is equivalent to saying that our domain in label space is just the lower half plane, which is much simpler to work with both theoretically and numerically. This is in contrast to in the Eulerian frame where the domain is bounded above by the free surface $\eta(x,t)$, which is itself a dependent variable of the system and not known a priori. Thus our surface boundary condition, equivalent to the dynamic boundary condition in Eulerian coordinates, simply states that pressure must vanish up to a constant at the surface, i.e.
@@
which is just another way of saying that the wave is unforced. We examine what happens when this condition is relaxed in a later section. The bottom boundary condition states that the vertical velocity must vanish as we tend towards the infinitely deep impermeable bottom.
As a final point, the fact that the domain in label space is time independent also means that all points initially within the domain remain there. This is in contrast to the Eulerian frame, where certain points, such as those with $y=0$, are outside of the fluid part of the time, and therefore taking temporal averages at these points becomes ill-defined.
@@ 3 Drift in relation to vorticity, momentum and energy
Up to this point we have neglected to mention the vorticity of these waves. While it has been long known that there exists an exact solution to the above system in which particles undergo purely circular trajectories, these Gerstner (1802) waves have a nonvanishing vorticity (Lamb, 1932, Art. 251). Irrotational waves are physically desirable since surface waves are assumed to be generated from an irrotational state of rest by the pressure gradient force, which as a conservative body force which cannot alter vorticity (Phillips, 1977). We can compute the vorticity in Lagrangian coordinates by a simple mapping between reference frames
@@
$\frac{{\rm d}q}{{\rm d}t}=\frac{\partial q}{\partial t}=0\,.$ (11)
Thus the vorticity in two-dimensional inviscid flow acts like a conserved ‘charge’ for particles, analogous to the electric charge in electromagnetism (see Salmon, 2014). This analogy is conceptually useful as well, as both electric charge and vorticity, even when vanishing, profoundly affect the motion of matter.
@@ 3.1 Series expansions
We consider permanent progressive and monochromatic waves and expand our trajectories in a series as
$x=\alpha+U(\beta)t+\sum_{n=1}^{\infty}x_{n}(\beta)\sin(\theta_{n})\,,\qquad y=\beta+y_{0}(\beta)+\sum_{n=1}^{\infty}y_{n}(\beta)\cos(\theta_{n})\,,$ (12)
with $\theta_{n}=nk(a-(c-U(\beta))t)$, $k$ the wavenumber, $c$ the phase speed, $U(\beta)$ the explicit mean Lagrangian drift, and $y_{0}(\beta)$ the mean water level, a parameter explored further in section 3.3. Note that the intrinsic frequency, Doppler-shifted by $U(\beta)$, is needed to remove secular terms at higher orders (see Clamond, 2007 on Buldakov et al., 2006). Inserting these expansions into the irrotational condition (10) and taking the time averaged component yields the kinematic constraint
@@
a result first found by Pizzo et al. (2023) which shows what form the drift must take to maintain irrotational flow. While this relation is purely kinematic, it shows that, given an expansion of the form (12), and so long as a wave is present (e.g. $x_{n}$,$y_{n}\neq 0$ for some $n$) there must be a positive definite mean Lagrangian drift $U(\beta)$ for the flow to stay irrotational. While this line of reasoning explains why a sheared mean flow is needed if an irrotational wave is present, we still lack a physical mechanism for its origin. To that end, we next turn to an investigation of wave dynamics.

### v3

# Momentum, energy and vorticity balances in deep-water surface gravity waves

@@ Abstract
The particle trajectories in irrotational, incompressible and inviscid deep-water surface gravity waves are open, leading to a net drift in the direction of wave propagation commonly referred to as the Stokes Drift, which is responsible for catalysing surface wave-induced mixing in the ocean and transporting marine debris. A balance between phase-averaged momentum density, kinetic energy density and vorticity for irrotational, monochromatic and spatially periodic two-dimensional water waves is derived by working directly within the Lagrangian reference frame, which tracks particle trajectories as a function of their labels and time. This balance should be expected as all three of these quantities are conserved following particles in this system. Vorticity in particular is always conserved along particles in two-dimensional inviscid flow, and as such even in its absence it is the value of the vorticity which fundamentally sets the drift, which in the Lagrangian frame is identified as the phase-averaged momentum density of the system. A relationship between the drift and the geometric mean water level of particles is found at the surface which highlights connections between geometry and dynamics. Finally, an example of an initially quiescent fluid driven by a wavelike pressure disturbance is considered, showing how the net momentum and energy from the surface pressure disturbance transfer to the wave field, recognizing the source of the mean Lagrangian drift as the net momentum required to generate an irrotational surface wave by any conservative force.
@@ 1 Introduction
Deep water surface gravity waves are ubiquitous in the global oceans, and affect the transport of heat, momentum and mass both along and across the air-sea interface (van Sebille, 2020; Melville, 1996; Deike, 2022). One crucial property of irrotational deep water waves is that the particle trajectories are not closed leading to a net drift in the direction of wave propagation commonly referred to as the Stokes Drift (Stokes, 1847). Formally, the Stokes drift is defined as the difference between the mean Lagrangian and mean Eulerian currents,
@@
where $\bm{u}_{L}$ and $\bm{u}_{E}$ represent the Lagrangian and Eulerian currents respectively, and the overline indicates a time mean in each reference frame over a wave period. One often neglected issue with this definition is the validity of taking the difference between two quantities in different reference frames with different dependent variables and more importantly different definitions of averaging as the Lagrangian and Eulerian periods are not equal (Longuet-Higgins, 1986). To avoid this confusion, we will instead use the term ‘wave-induced mean Lagrangian drift’ to refer to the mean Lagrangian velocity of fluid particles over the Lagrangian wave period.
The wave-induced mean Lagrangian drift modulates upper ocean currents, affects the transport of buoyant pollutants, plankton and marine debris (DiBenedetto et al., 2018) and enhances vertical mixing via Langmuir circulation (Craik & Leibovich, 1976; Belcher, 2012; Wagner et al., 2023). There is also evidence that this drift, or mean Lagrangian momentum density, can help with the interpretation of many central geometric, kinematic and dynamic properties of surface waves (Pizzo et al., 2023). Despite the elapse of over 175 years since its discovery, there is still confusion regarding the origins and interpretation of the wave-induced mean Lagrangian flow for irrotational surface gravity waves. Most derivations, including that of Stokes (1847), calculate the magnitude and direction of the drift from an asymptotic integration of the kinematic condition relating the Eulerian and Lagrangian velocities, which simply states that at a fixed point in time and space, the Eulerian and Lagrangian velocities are equal
@@
since at a fixed time a particle’s location is coincident with a fixed point in space. Thus the particle trajectories within a wave, which are fundamentally Lagrangian quantities, are typically derived from the Eulerian velocity fields. When done in this way, the mean Lagrangian drift appears to simply fall out of the math, and physical explanations for its existence tend to come post-factum. Why, fundamentally, should progressive irrotational surface waves induce a mean motion of water? What sets its magnitude and direction? Lastly, how is the mean Lagrangian drift related to other quantities such as vorticity and energy density? Such questions are the primary aim of this paper.
In section 2, we introduce the governing equations and relevant conditions for solving irrotational, incompressible, spatially periodic and inviscid two-dimensional progressive deep-water surface gravity waves in the Lagrangian reference frame, which tracks the trajectories of individual fluid parcels as a function of labelling coordinates. In this frame, the wave-induced mean Lagrangian drift is explicitly written as the average velocity of fluid parcels, and is thus identified as the mean Lagrangian momentum density of the system, physically motivated when one recalls momentum being equivalent to mass flux. In section 3, through an investigation of these equations, we show how the mean momentum density, or equivalently the drift, is related to the vorticity and energy density of irrotational water waves. Despite the flow being completely irrotational, it is precisely this very strict condition of irrotationality that dynamically mandates a sheared mean Lagrangian drift, and thus we emphasize that even in irrotational flow it is the vorticity that sets the magnitude and direction of the drift. We then dynamically connect the mean momentum and kinetic energy densities, showing that all monochromatic, irrotational and progressive waves with nonzero kinetic energy require a nonzero mean Lagrangian drift. Finally, we highlight a connection between the mean potential and mean kinetic energy densities through the mean pressure using a Bernoulli equation in the Lagrangian frame first outlined by Pizzo et al. (2023).
To further explore the dynamic relationship between these variables, in section 4 we consider the momentum and energy budgets within the physically motivated Lagrangian reference frame. To show how the wave-induced mean Lagrangian drift emerges from an initially quiescent flow, we consider a simple example where a still surface is resonantly driven by a wavelike pressure forcing. Through the integral momentum budgets, we find that all the momentum transferred to the waves from the surface forcing goes into the mean momentum, or mean Lagrangian drift. The same is shown to be true for the total energy. Finally, while the main body of this work concerns irrotational flow due to its prevalence in the literature, one can turn to the appendix for a generalization of each major result for rotational flow.
@@ 2 The Lagrangian description of water waves
Lagrangian quantities track evolution following fixed fluid particles. Thus a complete two-dimensional Lagrangian description of a fluid domain requires calculating particle trajectories $\bm{x}(\alpha,\beta,\tau)$ for each fluid element as a function of continuous particle labelling coordinates $(\alpha,\beta)$ and time $\tau$. Note that we distinguish $\tau$ from $t$ to emphasize that the partial derivative with respect to $\tau$ holds particle labels fixed. We will equivalently indicate such derivatives with an overhead dot. The particle trajectories $\bm{x}(\alpha,\beta,\tau)$ represent a general time-dependent mapping between label space $(\alpha,\beta)$ and physical space $(x,y)$ which must be invertible because no two particles can occupy the same physical location at the same time: that is, the function
@@
must not be equal to zero anywhere in the domain. Here subscripts indicate partial derivatives. The determinant of the Jacobian matrix of this mapping, $\mathcal{J}$, also allows us to easily change variables of differentiation. For example, the two dimensional incompressiblity condition in the Eulerian frame is denoted as
@@
which can be mapped to the Lagrangian frame through the following series of steps
$0=u_{x}+v_{y}=\frac{\partial(u,y)}{\partial(x,y)}+\frac{\partial(x,v)}{\partial(x,y)}=\frac{1}{\mathcal{J}}\bigg{(}\frac{\partial(u,y)}{\partial(\alpha,\beta)}+\frac{\partial(x,v)}{\partial(\alpha,\beta)}\bigg{)}=\\ \frac{1}{\mathcal{J}}\bigg{(}\frac{\partial(\dot{x},y)}{\partial(\alpha,\beta)}+\frac{\partial(x,\dot{y})}{\partial(\alpha,\beta)}\bigg{)}=\frac{1}{\mathcal{J}}\frac{\partial}{\partial\tau}\frac{\partial(x,y)}{\partial(\alpha,\beta)}=\frac{1}{\mathcal{J}}\dot{\mathcal{J}}=0\,.$ (5)
Therefore incompressible flow requires that the $\mathcal{J}$ be time independent. We could have determined this condition without calculation by remembering that the $\mathcal{J}$ determines how infinitesimal areas are mapped between label space and physical space. A small collection of particles ${\rm d}\alpha\,{\rm d}\beta$ must enclose the same physical area $\mathcal{J}^{-1}{\rm d}x\,{\rm d}y$ for all time or else the flow would be allowed to compress.
@@
where $p$ represents the pressure, and $g$ the acceleration due to gravity. Note that in the Lagrangian frame the nonlinear terms arise in the pressure terms and not in the inertia terms, in contrast to the Eulerian reference frame.
While incompressibility provides a constraint on our mapping between label space and physical space, there is still tremendous freedom in how we label our particles; this is known as the particle relabelling symmetry, and represents a gauge freedom of fluid mechanics. The conserved quantity associated with this gauge freedom is the vorticity (Salmon, 1988). Just as in electromagnetism, this gauge can be conveniently chosen to simplify computations, but we leave it general for now.
Here, our physical system amounts to solving equations (6), (7) for variables $(x,y,p)$ as functions of $(\alpha,\beta,\tau)$, subject to the incompressibility condition (5) for a given labelling gauge choice. To close the system, we impose boundary conditions at the free surface and the bottom. As part of our labelling freedom, we label particles at the surface with $\beta=0$, which makes the evaluation of surface quantities straightforward. This is equivalent to saying that our domain in label space is just the lower half plane, which is much simpler to work with both theoretically and numerically. This is in contrast to in the Eulerian frame where the domain is bounded above by the free surface $\eta(x,t)$, which is itself a dependent variable of the system and not known a priori. Thus our surface boundary condition, equivalent to the dynamic boundary condition in Eulerian coordinates, simply states that pressure must vanish at the surface, i.e.
@@
which is just another way of saying that the wave is unforced. We examine what happens when this condition is relaxed in a later section. The bottom boundary condition states that the vertical velocity must vanish as we tend towards the infinitely deep impermeable bottom
$\dot{y}(\beta=-\infty)=0\,.$ (9)
As a final point, the fact that the domain in label space is time independent also means that all points initially within the domain remain there. This is in contrast to the Eulerian frame, where certain points, such as those with $y=0$, are outside of the fluid part of the time, and therefore taking temporal averages at these points becomes ill-defined. The implications of these Eulerian averages are discussed further below.
@@ 3 Drift in relation to vorticity, momentum and energy
Up to this point we have neglected to mention the vorticity of these waves. While it has been long known that there exists an exact solution to the above system in which particles undergo purely circular trajectories, these Gerstner (1802) waves have a nonvanishing vorticity (Lamb, 1932, Art. 251). When surface waves are generated from a state of rest by a (conservative) pressure gradient force, they are irrotational (Phillips, 1977). We can compute the vorticity in Lagrangian coordinates by a simple mapping between reference frames
@@
$\dot{q}=0\,.$ (12)
However, this does not extend to three dimensions, where vorticity is no longer materially conserved on particles (due to vortex tilting and stretching), but is instead conserved on one-dimensional vortex lines. In two-dimensions, these lines collapse to a point because they are assumed to extend indefinitely in a direction orthogonal to the plane. Here, we restrict ourselves to two dimensional wave fields with no variation in the transverse direction. Applications to finite extent wave packets in both two and three dimensions is under current investigation by the authors (see also Pizzo & Salmon, 2021).
@@ 3.1 Series expansions
We consider permanent, progressive, spatially periodic and monochromatic waves in two dimensions and expand our trajectories in a series following Clamond (2007); Pizzo et al. (2023) as
$x=\alpha+U(\beta)\tau+\sum_{n=1}^{\infty}x_{n}(\beta)\sin(\theta_{n})\,,\qquad y=\beta+y_{0}(\beta)+\sum_{n=1}^{\infty}y_{n}(\beta)\cos(\theta_{n})\,,$ (13)
with $\theta_{n}=nk({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha}-(c-U(\beta))\tau)$, $k$ the wavenumber, $c$ the phase speed, $U(\beta)$ the explicit mean Lagrangian drift, and $y_{0}(\beta)$ the mean water level, a parameter explored further in section 3.3. The justification for these expansions is as follows. First, we expand about a rest state $(x=\alpha,y=\beta)$. We then include an explicit steady mean Lagrangian drift in $x$ and a mean water level in $y$, which will be constrained by the equations of motion and boundary conditions. Due to the horizontal periodicity of the wave profiles, a Fourier series expansion is best suited to represent the wavelike orbital motion of the fluid particles. Because the waves are permanent in shape, the Fourier coefficients can depend only on the vertical label $\beta$. Note that the intrinsic frequency, Doppler-shifted by $U(\beta)$, is needed to remove secular terms at higher orders (see Clamond, 2007 discussing Buldakov et al., 2006). Inserting these expansions into the irrotational condition (11) and taking the time averaged component yields
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-U^{\prime}(\beta)+(c-U(\beta))\sum_{n=1}^{\infty}n^{2}k^{2}(x_{n}^{\prime}(\beta)x_{n}(\beta)+y_{n}^{\prime}(\beta)y_{n}(\beta))=0\,,$ (14)
which can be integrated as
@@
a result first found by Pizzo et al. (2023) which shows what form the drift must take to maintain irrotational flow. The integration constant is zero since we are considering the frame where the fluid velocity vanishes at depth. While this relation just comes from a dynamic constraint, it shows that, given an expansion of the form (13), and so long as a wave is present (e.g. $x_{n}$,$y_{n}\neq 0$ for some $n$) there must be a positive definite mean Lagrangian drift $U(\beta)$ for the flow to stay irrotational. While this line of reasoning explains why a sheared mean flow is needed if an irrotational wave is present, we still lack a physical mechanism for its origin. To that end, we next turn to an investigation of wave dynamics.